Find a number such that the sum of 16 and that number is 5 times that number.
Answer by macston(5194)

.
Let the number be X:
.
The sum of the number and 16:
X+16
.
Five times that number:
5 x X=5X
.
X+16=5X . Subtract X from each side.
16=4X . Divide each side by 4.
4=X
ANSWER: The number is 4.
.
CHECK:
X+16=5X
4+16=5(4)
20=20